### v2

# Cognitive Paradigms for Evaluating VLMs on Visual Reasoning Task

Abstract
Advancing machine visual reasoning requires a deeper understanding of how Vision-Language Models (VLMs) process and interpret complex visual patterns. This work introduces a novel, cognitively-inspired evaluation framework to systematically analyze VLM reasoning on natural image-based Bongard Problems. We propose three structured paradigms—Direct Visual Rule Learning, Deductive Rule Learning, and Componential Analysis—designed to progressively enforce step-wise reasoning and disentangle the interplay between perception and reasoning. Our evaluation shows that advanced, closed-source VLMs (GPT-4o, Gemini 2.0) achieve near-superhuman performance, particularly when provided with high-quality image descriptions, while open-source models exhibit a significant performance bottleneck due to deficiencies in perception. An ablation study further confirms that perception, rather than reasoning, is the primary limiting factor, as open-source models apply extracted rules effectively when given accurate descriptions. These findings underscore the critical role of robust multimodal perception in enhancing generalizable visual reasoning and highlight the importance of structured, step-wise reasoning paradigms for advancing machine intelligence.
Cognitive Paradigms for Evaluating VLMs on Visual Reasoning Task
1 Introduction
Human intelligence seamlessly integrates visual perception with language and higher-level cognition, enabling sophisticated reasoning about the visual world Kunda (2020). This ability, known as visual reasoning, supports tasks ranging from understanding spatial relationships to inferring object functions. Despite advancements in natural language processing (NLP) Zhao et al. (2023), achieving robust visual reasoning in machines, especially in complex, multimodal contexts, remains a significant challenge Cao et al. (2024).
Large Language Models (LLMs) have revolutionized NLP, demonstrating impressive capabilities in tasks like machine translation, text generation, and question answering Brown et al. (2020); Devlin (2018). Their emergent abilities in areas such as code generation Chen et al. (2021) and logical reasoning Wei et al. (2022a) suggest their potential for tackling complex visual reasoning tasks. However, effectively bridging language and vision to achieve human-like visual reasoning requires careful consideration of how these models process and interpret visual information.
Visual reasoning challenges span a spectrum from basic object recognition Lin et al. (2014) to abstract inference Barrett et al. (2018); Zhang et al. (2019); Teney et al. (2020); Zerroug et al. (2022). One such challenge – Bongard Problems (BPs) Bongard (1968), involving abstract rules distinguishing two image sets, exemplify the challenges of integrating perceptual data with higher-order cognition. They require both perceptual and conceptual reasoning, making them a valuable testbed for evaluating visual reasoning capabilities.
Vision-Language Models (VLMs), such as Phi-1.5 Javaheripi et al. (2023), Llama-vision Meta (2024), Gemini 2.0 Google (2024), and GPT-4o OpenAI (2024), offer a promising avenue for addressing BPs. By leveraging multimodal training data, these models combine visual perception with higher-order cognitive processes. However, the mechanisms by which VLMs achieve visual reasoning, and how these mechanisms align with human cognitive strategies, remain an open question.
This paper investigates the visual reasoning abilities of VLMs on natural image-based Bongard Problems using the Bongard Openworld Dataset Wu et al. (2024). Inspired by human problem-solving strategies, we introduce three distinct paradigms—direct visual rule learning, deductive rule learning, and componential analysis—designed to dissect the interplay between perception and reasoning in VLMs. Our key contribution is the development and application of these cognitively-inspired paradigms, which allow us to isolate and analyze the distinct roles of perception and reasoning in VLM performance. We demonstrate that advanced VLMs can achieve near- and even super-human performance on this challenging task, but also reveal critical limitations in the perception modules of open-source VLMs. These findings offer valuable insights into the computational underpinnings of visual reasoning and provide crucial directions for future research.
Research Contributions:
  1. 1.
    Three distinct, cognitively-inspired paradigms for evaluating VLM’s visual reasoning ability: Direct Visual Rule Learning, Deductive Rule Learning, and Componential Analysis, isolating the roles of perception and reasoning.
  2. 2.
    Demonstration of near- and super-human performance by VLMs, particularly with high-quality image descriptions.
  3. 3.
    Identification of a critical bottleneck in open-source VLMs: the perception module (image description generation), which significantly limits their reasoning capabilities.
  4. 4.
    Highlighting the crucial role of robust image descriptions for complex visual reasoning and the need for future research to focus on improving the perception components of open-source VLMs.
2 Related Work
Evaluating Vision-Language Models (VLMs) requires diverse benchmarks targeting various aspects of visual understanding and reasoning. Traditional benchmarks focused on single-image tasks like image captioning Lin et al. (2014), OCR Mishra et al. (2019); Singh et al. (2021), and VQA Antol et al. (2015); Zhang et al. (2024); Lu et al. (2022); Gurari et al. (2018). However, the field has increasingly shifted towards more complex scenarios, including multi-image comprehension. This shift has led to the development of interleaved image-text corpora like MMC4 Zhu et al. (2024) and OBELICS Laurençon et al. (2024); Lin et al. (2024), along with dedicated benchmarks specifically designed for multi-image understanding, such as Q-Bench Wu et al. (2023), Memontos Wang et al. (2024), and DEMON Li et al. (2023). Benchmarks like MMIU Meng et al. (2024) further broaden the scope by incorporating diverse datasets. While these benchmarks represent significant progress, they often fall short in capturing the intricacies of abstract visual reasoning, particularly those devoid of linguistic input.
Purely visual reasoning tasks, where language is absent, are crucial for isolating and evaluating core visual understanding. Bongard problems Bongard (1968) provide an excellent example of such tasks, focusing on abstract visual concepts and requiring multi-image relational reasoning. Datasets like Bongard-HOI Jiang et al. (2022) and Bongard-OpenWorld Wu et al. (2024) extend these challenges to human-object interactions and real-world scenarios, respectively, pushing models to demonstrate robust few-shot reasoning capabilities.
This focus on purely visual reasoning is not unique to Bongard problems. Other notable benchmarks in this category include SVRT Fleuret et al. (2011), CVR Zerroug et al. (2022), Raven’s Progressive Matrices (RPM) Barrett et al. (2018); Zhang et al. (2019), V-PROM Teney et al. (2020), and ARC Chollet (2019). These benchmarks collectively assess a wide range of visual reasoning abilities, including spatial, compositional, and abstract reasoning, offering valuable insights into the visual cognition of AI models.
Our work is also informed by research in human cognition, which has identified key learning strategies relevant to visual reasoning. We consider three key learning strategies that are relevant to visual reasoning: direct visual rule learning – integrating all information simultaneously, akin to similarity-based strategies von Helversen et al. (2014); Johnson-Laird (2010), deductive reasoning – prioritizing rule extraction and application Knauff and Wolf (2010); Helie et al. (2015); Rips (1994), and componential analysis – breaking down problems into individual components Wu and Molnár (2022); Gluck et al. (2008); Hutchinson (1992). These strategies inform the design and interpretation of our visual reasoning benchmarks, enabling us to assess how well VLMs emulate different facets of human cognition.
Finally, our work has connections to the growing body of research on Chain-of-Thought (CoT) prompting Kojima et al. (2022); Wei et al. (2022b); Huang and Chang (2023). While our focus is on visual reasoning and human cognitive strategies, CoT prompting, which encourages models to generate intermediate reasoning steps, shares a similar goal of improving model reasoning abilities. Recent work has explored applying CoT to multimodal tasks Zhang et al. (2023); Chen et al. (2024); Shao et al. (2024), demonstrating its potential for enhancing performance. Our work, by explicitly separating to visual modality alone focusing on perception and reasoning and analyzing the models’ ability to extract and apply rules, provides a complementary perspective on understanding and improving visual reasoning in VLMs.
3 Models
To evaluate our approach comprehensively, we employed a diverse range of Vision-Language Models (VLMs) that include both closed-source and open-source architectures. This diversity allowed us to assess the strengths and limitations of different model types and scales. Our selection encompasses state-of-the-art multimodal models, including GPT-4o OpenAI (2024), Gemini 2.0-flash-exp, henceforth Gemini 2.0 Google (2024), Pixtral-12B Agrawal et al. (2024), Llama-Vision-3.2 (11B, and 90B) Meta (2024), LLaVA (7B, 13B and 34B) Liu et al. (2023), and LLaVA built on Llama 3 (8B) XTuner (2025). By leveraging this variety, we systematically analyzed model performance across different architectural designs and parameter scales, offering a holistic view.
For the evaluation purpose, two metrics are used: classification accuracy and semantic similarity. Further technical details, including model architectures, evaluation and individual summaries, are elaborated in Appendix A.4.
4 Dataset
We evaluate our approach on the Bongard OpenWorld dataset Wu et al. (2024), a benchmark designed to assess few-shot visual reasoning abilities. The full dataset contains 1001 cases, from which we have chosen a sample of equally balanced 500 test samples for our experiments (further details can be found in Appendix A.3). Each case consists of six positive and six negative examples and a query image. These cases are governed by an underlying “commonsense” rule, with the rules themselves categorized into distinct visual concepts. Figure A.1 illustrates an example task where the identifying rule is to detect a group wedding photo.
5 Experiments
This paper investigates VLMs’ ability to solve natural image-based Bongard Problems. We evaluate three paradigms: direct visual rule learning (processing all information together), deductive rule learning (deriving rules first), and componential analysis (examining components separately). For each paradigm, the VLMs were prompted to provide four distinct outputs: (1) an overall analysis of the visual features and their relationships, (2) the derived rule that distinguishes the positive and negative categories, (3) a description of the query image, and (4) the final classification category for the query image. This consistent, structured output format facilitated a detailed examination of the models’ reasoning processes across all paradigms.
Figure 1: A Multi-Stage Evaluation Framework for Visual Reasoning. Direct Visual Rule Learning: Positive and negative examples, along with a query image, are processed concurrently to identify the distinguishing rule and categorize. Deductive Rule Learning: A rule is extracted from the positive and negative examples and then applied to classify the query image. Componential Analysis: Textual descriptions of the images are generated, and rule extraction and classification are performed based solely on these textual representations.
5.1 Direct Visual Rule Learning
We begin by exploring a holistic approach to visual reasoning. In this paradigm, the Vision-Language Model (VLM) is presented with all images relevant to a given Bongard problem – the positive examples, the negative examples, and the query image – simultaneously. This approach aims to mimic human problem-solving strategies where all available information is considered concurrently. The prompt instructs the model to: (1) provide the overall analysis, (2) identify a rule that distinguishes the positive examples from the negative examples; (3) analyze the query image; (4) classify the query image as either positive or negative based on the identified rule The prompt is detailed in Appendix A.5.1.
5.2 Deductive Rule Learning
A core aspect of human intelligence is the ability to deduce general rules from specific examples and then apply those rules to new situations. This paradigm emulates human deductive reasoning, involving two key steps: first, inferring a general rule from a set of positive and negative examples; second, applying this derived rule to classify a query image.
To investigate this capability in VLMs, we designed a two-stage prompting strategy. In the first stage, the model was presented with six positive examples and negative examples. The prompt, detailed in Appendix A.5.2.1, instructed the model to identify a distinguishing rule that separates the two categories. Critically, the model was asked to provide a concise summary of this identified rule, limited to a maximum of 20 words. This constraint encouraged the model to focus on the most salient features distinguishing the categories.
In the second stage, the model was provided with the previously generated rule summary and a query image. The prompt, detailed in Appendix A.5.2.2, instructed the model to analyze the query image in the context of the given rule and classify it as either positive or negative. The model was also asked to provide a brief analysis of the query image and restate the rule it was applying. This allowed us to examine not only the classification decision but also the model’s reasoning process.
5.3 Componential Analysis
While deductive rule learning operates at the level of abstract rules, componential analysis takes a more granular approach. This method decomposes each image into its constituent visual elements and reasons based on these components, represented textually. This approach mirrors how humans often break down complex visual scenes into simpler parts for analysis. The process unfolds in two distinct stages: (1) detailed image description generation and (2) rule derivation and application.
The componential analysis experiments aim to evaluate the ability of VLMs to reason based on structured, textual descriptions of visual content. We employed a two-stage prompting strategy to simulate this compositional reasoning process. In the first stage, detailed descriptions of each image (positive, negative, and query) were generated using the prompt detailed in Appendix A.5.3.1. This prompt instructs the model to analyze an image and provide a structured textual description covering aspects such as scene, objects (living and inanimate), activities, contextual elements, visual patterns, emotional undertones, textual information, and a summary.
In the second stage, the model was provided with the descriptions generated in the first stage, along with explicit “textual” information about which images correspond to positive examples and which correspond to negative examples, as well as the description of the query image. The prompt, detailed in Appendix A.5.3.2, then instructed the model to derive a rule based on the descriptions and apply this rule to classify the query image – described in textual form. This stage combines the processes of rule derivation and application, relying solely on the textual representations of the visual content. This isolates the models’ ability to reason about visual concepts based on their textual representations.
6 Results
6.1 Direct Visual Rule Learning
Direct visual rule learning necessitates models with multi-image and multimodal capabilities. Among the models tested, only Gemini 2.0 and GPT-4o possess these features.
Table 1 presents the overall accuracy, as well as the accuracy on positive and negative samples separately. As shown, Gemini 2.0 achieves a slightly higher overall accuracy of 82.2% compared to GPT-4o’s 80.0%. These results suggest that while GPT-4o may excel at recognizing patterns within the positive examples, its lower accuracy on negative samples indicates challenges in distinguishing subtle counter examples. On the other hand, Gemini 2.0’s balanced performance reflects greater robustness in addressing the diverse reasoning demands of direct visual rule learning tasks.
6.2 Deductive Rule Learning
The deductive rule learning paradigm, with its two-stage process of rule extraction and application, offers a more structured approach to visual reasoning compared to the direct visual rule learning. This structured approach allows us to evaluate the model’s ability to not only identify patterns but also articulate them as explicit rules, providing insights into its deductive reasoning capabilities. We conducted these experiments using Gemini 2.0 and GPT-4o. Table 1 summarizes the overall accuracy, as well as the accuracy on positive and negative samples separately.
As shown in Table 1, GPT-4o achieves slightly higher overall accuracy than Gemini 2.0, suggesting greater effectiveness in combining rule extraction and application within this paradigm. While both models demonstrate relatively balanced performance across positive and negative samples, subtle differences emerge. Similar to the previous experiment, GPT-4o exhibits a stronger ability to classify positive samples, and Gemini 2.0 achieves more consistent performance across both categories, suggesting greater robustness across diverse reasoning scenarios.
Comparing these results to those obtained using the direct visual rule learning (Table 1), we can see a clear improvement in overall accuracy for both models. This indicates that the two-stage deductive approach enhances the models’ ability to handle the complexities of both positive and negative test samples, surpassing the performance seen in the holistic approach. This improvement likely stems from the explicit separation of the rule extraction and application processes, allowing the models to focus on each task individually. This suggests that explicitly structuring the reasoning process, even within purely visual tasks, benefits VLMs.
Model Direct Visual Rule Learning Deductive Rule Learning Componential Analysis
neg pos Overall neg pos Overall neg pos Overall
GPT-4o 66.4 93.6 80.0 82.8 93.2 88.0 92.8 92.8 92.8
Gemini 2.0 80.5 83.9 82.2 85.7 87.9 86.8 90.8 96.4 93.6
Pixtral-12B - - - - - - 59.2 48.4 53.8
Llava-Llama3-8B - - - - - - 93.0 13.1 53.2
Llama-vision-11B - - - - - - 42.8 64.0 53.4
Llava 7B - - - - - - 65.2 67.2 66.2
Human - - - - - - - - 91.0
Table 1: Classification accuracy across different Analysis Paradigms and Human performance Wu et al. (2024). neg and pos represents the classification accuracy for negative and positive query images.
6.3 Componential Analysis
The componential analysis approach evaluates VLMs by decomposing the complex multi-image reasoning task into manageable single-image description tasks. This method separates the perception component (image analysis and description) from the reasoning component (rule derivation and application), relying solely on textual representations for the latter. This isolation of perception and reasoning allows for a deeper understanding of their respective contributions to VLM performance.
Table 1 presents the performance metrics for six VLMs: Pixtral-12B, Llava-Llama3-8B, Llama-Vision-11B, Llava 7B, GPT-4o, and Gemini 2.0. The results reveal significant performance variations, highlighting the crucial role of accurate image descriptions in complex visual reasoning. GPT-4o and Gemini 2.0 achieve near-optimal accuracies (92.8% and 93.6%, respectively), surpassing human-level performance i.e. 91% Wu et al. (2024). This demonstrates the potential of advanced VLMs when provided with high-quality, structured descriptions. Their strong performance suggests that these models excel at leveraging textual information for rule derivation and application—the reasoning component.
In contrast, the open-source models exhibit greater performance variability. While Pixtral-12B and Llava 7B achieve moderate, balanced performance, Llava-Llama3-8B and Llama-Vision-11B show a striking bias: one performs well on negative samples but poorly on positive samples, while the other exhibits the opposite trend. This suggests a significant challenge in generalization for these models, likely stemming from inconsistencies or inaccuracies in their generated descriptions—the perception component.
The stark difference in performance between the closed- and the open-source models underscores the importance of robust image description generation. The near- or super-human performance of GPT-4o and Gemini 2.0 highlights the potential of VLMs when both perception and reasoning components are strong. The challenges faced by the open-source models, suggest that future research should prioritize improving the perception modules of these models to achieve human-like reasoning. Further investigation is needed to understand the specific factors contributing to the open-source models’ difficulties in generating consistent and accurate descriptions.
7 Ablation
7.1 Rule-Based Evaluation
The Deductive Rule Learning paradigm (Section 6.2) combines rule extraction and application. To isolate the latter, we designed an ablation study focusing solely on rule application, thus disentangling perception from reasoning. This allows us to directly assess the models’ reasoning capabilities independent of their perception modules.
Specifically, we used the rule summaries generated by GPT-4o during the rule extraction phase. For each Bongard problem, we provided the open-source models with two inputs: 1) the corresponding rule summary, and 2) the query image. The models were instructed to classify the query image based on the provided rule extracted using GPT-4o. This isolates the models’ ability to understand and apply a given rule, independent of their capacity to extract it.
Table 2 presents the results. Pixtral-12B achieved the highest overall accuracy (83.8%), demonstrating strong rule comprehension and application. Its balanced performance suggests robust rule application across contexts, contrasting with some other models exhibiting a bias towards confirming positive matches. While some achieve high positive accuracy, they struggle with negative cases, suggesting difficulty identifying when a rule does not apply. Pixtral-12B’s balanced performance suggests greater adeptness, while models like Llama-Vision-xx and LLaVA-xx may over-rely on confirmatory evidence.
Comparing the performance of Pixtral-12B and the Llava models with their componential analysis results reveals a crucial insight. Pixtral-12B maintains relatively consistent performance across both evaluations. However, the Llava models’ accuracy drops substantially in the componential analysis compared to the rule-based evaluation. This performance gap strongly indicates that the Llava models’ primary bottleneck lies in the perception stage – the ability to generate accurate and comprehensive image descriptions – rather than in their reasoning capabilities (rule application), which are more effectively assessed when pre-extracted rules are provided.
By isolating rule application, this ablation provides insights into the open-source models’ strengths and limitations in reasoning. Combined with the Deductive Rule Learning results, this analysis disentangles factors contributing to overall performance, distinguishing between deficits in rule extraction (a proxy for perception) and application (pure reasoning).
Model Accuracy (%)
Overall neg pos
LLaVA-7B 66.2 65.2 67.2
Llama-vision-11B 68.2 38.0 98.4
Pixtral-12B 83.8 75.2 92.4
LLaVA-13B 70.0 58.8 81.2
LLaVA-34B 74.8 52.8 96.8
Llama-vision-90B 74.2 59.6 88.8
Table 2: Rule-Based Evaluation Results showing overall classification accuracy along with categorical accuracy for positive (pos) and negative (neg) samples.
7.2 Analysis per Commonsense Category
To better understand the comparative capabilities of GPT-4o and Gemini 2.0, we analyzed their performance across ten commonsense reasoning categories – provided in the Bongard Openworld dataset (details in Appendix A.3.3). This analysis highlights the models’ strengths and limitations in handling diverse commonsense reasoning categories (refer to Table A.3). Both models demonstrated robust performance across most categories indicating models’ capability to process high-level conceptual patterns effectively.
Differences in performance were observed in categories requiring more nuanced or contextual understanding, such as “Taste / Nutrition / Food” and “Human-Object Interaction.” This suggests that while both models are adept at general reasoning tasks, their architecture and training data influence their performance on specific types of visual information. Notably, Gemini 2.0 showed strengths in scenarios involving fine-grained object attributes, while GPT-4o performed better in tasks demanding contextual interpretation.
7.3 Semantic Similarity Analysis
To further analyze the reasoning process of the best-performing models, GPT-4o and Gemini 2.0, we conducted a semantic similarity analysis. This analysis aims to quantify the alignment between the perception (image description) and the reasoning component (identified rule). We hypothesized that for correctly classified positive examples, the similarity between the description and rule should be high, reflecting a strong connection between perception and the derived rule. Conversely, for negative examples, the similarity should be lower, as the description should not align with the rule.
Using cosine similarity of sentence embeddings generated using OpenAI embeddings, we computed the mean and standard deviation of the semantic similarity scores for both models and categories. Both models exhibit higher similarity (see Table A.4) for positive examples compared to negative ones, confirming that correct classifications align with the identified rule. However, the small difference in similarity ($\sim$ 6%) suggests that the negative samples were designed to be closer to positive samples to make them challenging. This highlights the subtle perceptual distinctions required for correct classification in these challenging cases.
7.4 Analysis of Incorrectly Classified Samples
Here we analyzed the misclassified samples from GPT-4o and Gemini 2.0 during the Componential Analysis experiments to identify common limitations. The total 15 common test samples were incorrectly identified by both the VLMs. The errors revealed several recurring issues (more details in Table LABEL:tab:tasks). We found eleven cases of errors in rule extraction; more specifically, four cases of being too general, one case of a missed property, five cases of the missing required object, two cases of focusing on a wrong property. We also found three cases of weak similarity-based reasoning disregarding some parts of the rule. These findings highlight that while the models excel in many areas, they still face challenges in handling complex visual details and reasoning, suggesting areas for improvement in VLMs’ robustness and generalization.
8 Discussion
This study investigates how VLMs tackle Bongard Problems (BPs), offering insights into their visual reasoning mechanisms and alignment with human cognitive strategies. We introduced three evaluation paradigms, each mirroring a distinct human problem-solving approach. These paradigms allow us to evaluate VLM performance and dissect the contributions of different cognitive processes to their visual reasoning abilities.
Human scene understanding is an active, interactive process Zelinsky (2013); Malcolm et al. (2016); Gibson (2015). This interactionist perspective is key to understanding how both humans and VLMs approach BPs. Rather than passively receiving sensory input, both benefit from actively considering the relationships between images. Our Direct Visual Rule Learning paradigm directly probes this, requiring simultaneous processing of all images, mirroring holistic human problem-solving where a rapid “gist” perception guides further analysis Biederman (1987); Li et al. (2002). This initial holistic understanding is then typically refined through focused attention Posner (1980). Our Deductive Rule Learning paradigm simulates this by prompting VLMs to first learn the underlying rule before evaluating query images (more in Appendix A.2). This mimics human deductive reasoning where a hypothesis is formed and then tested.
Humans also frequently employ analytical strategies, breaking down complex problems. Our Componential Analysis paradigm reflects this, with VLMs analyzing positive and negative examples separately before comparison. This aligns with dual-process theories of cognition, where intuitive System 1 - perception, precedes deliberate System 2 - reasoning Kahneman (2011); Marr (2010). Critically, this paradigm allows us to isolate the perception component (image description) from the reasoning component (rule derivation and application), revealing the distinct roles these play in VLM performance.
Importantly, our three paradigms can be seen as progressively more detailed instantiations of a Chain-of-Thought (CoT) approach (more details in Appendix A.6.3). Direct Visual Rule Learning implicitly requires holistic image relationship consideration, Deductive Rule Learning adds explicit rule extraction and Componential Analysis represents the most granular CoT, decomposing images into separate descriptions. This progressive increase in detail is mirrored by increased performance, suggesting that structuring visual reasoning benefits VLMs. Unlike typical CoT prompting, our approach uses structured visual and textual representations in multiple stages, offering a way to induce step-by-step reasoning.
Our results show that VLMs can surpass human performance on BPs given high-quality, structured descriptions, notably in the Componential Analysis paradigm. This suggests that mitigating perceptual challenges allows VLMs to effectively use textual representations for sophisticated reasoning. However, the substantial performance gap between open- and closed-source models reveals critical limitations, especially in the perception component. The stark contrast between rule-based evaluation (bypassing perception) and Componential Analysis (requiring it) underscores this bottleneck. For instance, Llava models exhibited a significant performance drop when generating descriptions compared to using pre-extracted rules.
9 Limitations
While this work offers valuable insights, it’s crucial to acknowledge its limitations. Our evaluation relies on the Bongard OpenWorld dataset, which may not fully capture the diversity of real-world visual reasoning scenarios. Exploring more diverse datasets is essential for assessing generalizability. The Componential Analysis paradigm, while insightful, is inherently tied to the quality of generated image descriptions, which are often imperfect, especially for open-source VLMs. Furthermore, the computational demands of state-of-the-art models raise concerns about scalability and accessibility. Finally, our quantitative metrics may not fully capture the nuances of model reasoning.
Future work should prioritize improving open-source VLMs’ perception modules, exploring more diverse datasets, and developing richer evaluation metrics to provide deeper insights into model reasoning. Integrating VLMs and LLMs to leverage their complementary strengths is a promising direction for achieving more human-like reasoning.
10 Conclusion
This work presents a cognitively-inspired framework for evaluating the visual reasoning capabilities of Vision-Language Models (VLMs) on natural image-based Bongard Problems. By systematically dissecting the interplay between perception and reasoning, we introduce three structured paradigms—Direct Visual Rule Learning, Deductive Rule Learning, and Componential Analysis—that progressively enforce step-wise reasoning. Our findings reveal that advanced, closed-source VLMs (GPT-4o, Gemini 2.0) achieve near-superhuman performance, particularly when provided with high-quality image descriptions, whereas open-source models exhibit a significant performance gap due to deficiencies in their perception modules. The ablation study further highlights that perception, rather than reasoning, is the primary bottleneck, as open-source models can apply extracted rules effectively when given accurate descriptions. These results underscore the critical role of robust multimodal perception in advancing generalizable and interpretable visual reasoning models. Addressing this challenge requires enhancing visual representations, improving multimodal alignment, and designing architectures that enforce structured, step-wise reasoning akin to Chain-of-Thought prompting. Additionally, integrating cognitive principles from human problem-solving strategies could provide a pathway toward developing more robust, transparent, and human-aligned visual reasoning systems.
Acknowledgments
This work was supported by the Estonian Centre of Excellence in AI [grant number TK213U8]
Appendix
Appendix A.1 Broader Relevance
This study offers insights with broader implications for the development of more robust and human-like AI systems. Our cognitive inspired evaluation paradigms — Direct Visual Rule Learning, Deductive Rule Learning, and Componential Analysis—provide valuable tools for assessing and understanding the strengths and limitations of Vision-Language Models (VLMs) on complex visual reasoning tasks. The insights gained from applying these paradigms to the Bongard Open-World Dataset extend beyond this specific task, contributing to the development of VLMs capable of advanced reasoning in real-world applications.
Our key finding—that the perception module (image description) is a critical bottleneck for open-source VLMs—has significant implications for future research. Improving the ability of these models to generate accurate and comprehensive image descriptions is crucial for bridging the performance gap with closed-source models and unlocking their full potential for complex visual reasoning.
The demonstration that advanced VLMs (GPT-4o and Gemini 2.0) can achieve near- or even super-human performance on BPs, particularly when provided with high-quality descriptions, underscores the potential of these models for tasks requiring sophisticated visual understanding. This reinforces the importance of continued research into multimodal architectures that effectively integrate robust perception with advanced reasoning capabilities.
Finally, our comparative evaluation of open- and closed-source models contributes to the broader discussion about AI accessibility and transparency. By highlighting the performance disparities, we identify specific areas for improvement in open-source models, promoting the development of more accessible and capable AI systems. This work provides a foundation for future research aimed at bridging the gap between human and machine visual reasoning, paving the way for more robust and reliable AI in diverse real-world applications.
Appendix A.2 Attention and Memory in Visual Reasoning
While our study primarily focuses on the interplay between perception and reasoning, the roles of attention and memory are also implicitly present in our paradigms. The Direct Visual Rule Learning paradigm, by requiring simultaneous processing of all images, likely engages VLMs’ “visual attention” mechanisms to identify salient features and relationships across the image set. This aligns with research on human attention, which suggests that holistic scene processing relies on rapid perception and attentional guidance Biederman (1987); Li et al. (2002). These mechanisms allow VLMs to selectively focus on relevant image features, mirroring human gaze patterns and attentional shifts. Similarly, the Deductive Rule Learning paradigm relies on the model’s ability to “memorize” the extracted rule for later application to query images, a process that likely involves both working memory Baddeley (2012) and long-term memory storage Squire (1992). Memory mechanisms are essential for encoding visual information, maintaining it during rule formulation, and retrieving it for comparison. This is reflected in the VLMs’ ability to store learned rules (long-term memory) and maintain representations of the images during analysis (working memory). Although we did not directly manipulate or measure these cognitive processes, their involvement is undeniable, and future research could explore these aspects more explicitly, perhaps by analyzing attention maps Bahdanau et al. (2016); Vaswani et al. (2017) or investigating the memory representations learned by VLMs during rule extraction. The interplay between attention and memory, crucial for both human and artificial intelligence, provides valuable insights into the computational underpinnings of visual reasoning Vaishnav and Serre (2023).
Appendix A.3 Dataset
A.3.1 Bongard OpenWorld Dataset
We utilize a subset of 500 test cases from the Bongard OpenWorld dataset Wu et al. (2024), a benchmark for evaluating few-shot reasoning in machine vision. Each of the 1001 Bongard samples in this dataset contains 7 positive and 7 negative images, distinguished by an underlying "commonsense" rule represented by numerical values (0-9). To create our evaluation set, we followed a specific sampling strategy. Each test case in our subset consists of 6 positive images, 6 negative images, and 1 query image. From each of the original 1001 Bongard samples, we can generate up to 14 unique, balanced test cases using combinations. For our evaluation, we selected the first 250 Bongard samples from the full dataset. From each of these 250 samples, we selected one positive and one negative test case, resulting in a total of 500 test cases for our analysis. The specific samples used in our analysis will be made publicly available upon acceptance of this work. These rules encode different visual and spatial concepts, as detailed below.
Figure A.1: Illustration of a Bongard OpenWorld dataset showing three positive and negative samples along with a query image. Distinguishing rule in this example is A group photo at a wedding reception. Based on the rule, query image belongs to negative category.
A.3.2 Commonsense Value Categories
The Bongard OpenWorld dataset categorizes the underlying rules into several visual concept categories, some of which are explicitly related to commonsense reasoning. Table A.1 provides a summary of these categories and example concepts.
Concept Category ID Example
Anything else 0 Animals are running.
Human-Object Interaction (HOI) 1 A person playing the guitar.
Taste / Nutrition / Food 2 A plate of high-calorie food.
Color / Material / Shape 3 A wooden floor in the living room.
Functionality / Status / Affordance 4 An animal capable of flying in the tree.
And / Or / Not 5 A man without beard.
Factual Knowledge 6 A building in US capital.
Meta Class 7 Felidae animals.
Relationship 8 A bench near trees.
Unusual Observations 9 Refraction of light on a glass cup.
Table A.1: Commonsense ID Categories in Bongard OpenWorld
A.3.3 Commonsense Value Distribution in Our Subset
Table A.2 provides a detailed breakdown of the distribution of commonsense values within our 500 test-cases subset. As observed, the value ‘0’ (‘Anything else’, marked as commonsense) is significantly more prevalent than other values. This distribution should be considered when analyzing the model’s performance, as it may indicate a bias towards learning certain types of rules, particularly those related to the ‘Anything else’ category.
ID Count Percentage
0 365 73%
3 36 7.2%
9 26 5.2%
1 15 3%
2 14 2.8%
4 12 2.4%
5 10 2%
6 10 2%
8 8 1.6%
7 4 0.8%
Total 500 100%
Table A.2: Distribution of Commonsense ID in Bongard OpenWorld Dataset test cases used for experiments
A.3.4 Dataset Availability
The complete Bongard OpenWorld dataset is publicly available at Bongard-OpenWorld.github.io.
Appendix A.4 Model Details
This section provides comprehensive details on the Vision-Language Models (VLMs) utilized in our experiments.
A.4.1 Vision-Language Models (VLMs)
  • ChatGPT (GPT-4o with vision): OpenAI’s GPT-4o, accessed with vision capabilities via the openai library, processes both images and text seamlessly. It OpenAI (2024) excels in reasoning and instruction-following tasks, offering robust text generation and broad general knowledge. However, its closed-source nature and cost make it less accessible for extensive experimentation. Accessed version of the model is 2024-08-06.
  • Gemini (2.0-flash-exp): Google’s Gemini is a multimodal model Google (2024) emphasizing speed and computational efficiency in its “flash” version, accessed through the google.generativeai library. While it provides rapid inference capabilities, it is an experimental model with limited public documentation, restricting detailed architectural insights.
  • Pixtral (12B): Accessed via the mistralai library, Pixtral is a 12-billion-parameter model designed for vision-language tasks Agrawal et al. (2024). Its architecture and training details are not fully disclosed, but it shows strong performance in image-text reasoning. As a proprietary model, its access is constrained to specific APIs.
  • Llama-Vision (3.2, 11B, and 90B): Llama-Vision Meta (2024) combines a vision encoder with the Llama language model, with experiments conducted on 11B and 90B parameter variants to evaluate the impact of scaling. Accessed via the ollama library, its open-source availability enables detailed exploration, but the 90B version requires significant computational resources.
  • LLaVA (7B, 13B and 34B): LLaVA (Large Language and Vision Assistant) integrates a vision encoder with a large language model, fine-tuned for visual instruction-following tasks Liu et al. (2023). Accessed through the ollama library, it is particularly effective in visual question answering but demands substantial GPU memory.
  • LLaVA-Llama 3 (8B): This variant of LLaVA incorporates the Llama 3 architecture XTuner (2025). Accessed via the ollama library, it benefits from the advancements of Llama 3 Dubey et al. (2024), offering improved efficiency over larger models. However, its relative novelty means limited community feedback and tooling support.
A.4.2 Model Configurations and Hyperparameters
For closed-source models, we utilized API calls by installing the openai, mistralai (for Pixtral), and google.generativeai libraries, enabling access to ChatGPT (GPT-4o with vision) OpenAI (2024), Pixtral (12B) Agrawal et al. (2024), and Gemini 2.0 (flash-exp) Google (2024), respectively. The open-source models, on the other hand, were accessed using the ollama Ollama (2025) library.
  • Input Format: Bongard problems were presented by encoding images in base64 and embedding them within text prompts. For instance: “Here are two sets of images. The first set represents positive examples, and the second set represents negative examples of a concept. What is the rule that distinguishes the two sets? ”
  • Image Size: For OpenAI and Gemini 2.0 models, the default image sizes provided by their APIs were used. For all other models, the maximum image size was restricted to 1024 pixels to ensure uniformity.
  • Decoding Parameters: The temperature parameter was set to 0 to maintain deterministic outputs during evaluation.
  • Fine-tuning: No fine-tuning was performed; all models were evaluated in their pre-trained configurations.
  • Hardware: Experiments were conducted using NVIDIA GPUs, including 2080Ti, 3090, and 6000 Ada, ensuring adequate computational power for all model evaluations.
A.4.3 Evaluation Metrics
Classification Metric:
For evaluating the performance of the VLMs, we employ categorization accuracy on the binary classification task. This metric measures the percentage of test instances correctly classified as either positive or negative as present in the Bongard Problems.
Semantic Similarity:
We leverage a Semantic Similarity Evaluator to quantify the semantic relatedness between generated responses (e.g., question answers) and ground truth answers. This evaluator is inspired by the work of Risch et al. (2021), which explores semantic answer similarity for evaluating question answering models.
Appendix A.5 Model Prompts
A.5.1 Direct Visual Rule Learning
The prompt used for the Direct Visual Rule Learning paradigm is designed to elicit a holistic analysis of the provided images, encouraging the model to identify a distinguishing rule and apply it to the query image. The prompt emphasizes the distinction between positive ($cat\_2$) and negative ($cat\_1$) examples and guides the model to provide a structured output containing its analysis, the identified rule, details about the query image, and the final classification.
def visual_concept_test_prompt(m, n):
Generates a visual analysis prompt.
Args:
m (int): Number of positive samples.
n (int): Number of negative samples.
Returns:
str: The formatted prompt string.
return f"""
You are provided with {m + n + 1} images: the first {m} samples are ‘cat_2‘, the next {n} samples are ‘cat_1‘, and the last image is the ‘query image‘.
Analyze the common characteristics or patterns found in the ‘cat_2‘ samples (positive samples: following 1 common rule) that distinctly separate them from the ‘cat_1‘ samples (negative samples: it might not follow any possible rule).
Your task is to:
1. Determine the rule or criterion that distinguishes the ‘cat_2‘ samples from the ‘cat_1‘ ones.
2. Analyse the ‘query image‘ (last image).
3. Provide your conclusion for the ‘query image‘ if it can be categorized as either ‘cat_1‘ or ‘cat_2‘ based on the analysis and the rule.
Ensure that the output is clear, well-formatted, and free of unnecessary explanations.
Omit the ‘‘‘ tags at the beginning and end of the page. The format of your output should be as follows:
- **Analysis**: (Your analysis here)
- **Rule**: (The distinguishing rule here)
- **Query Image**: (Query image details)
- **Conclusion**: (cat_1 or cat_2)
A.5.2 Deductive Rule Learning
The Deductive Rule Learning paradigm employs a two-stage prompting strategy. The first stage focuses on rule extraction from positive and negative examples, while the second stage applies the extracted rule to classify a query image. The prompts for each stage are detailed below.
A.5.2.1 First-Stage Prompt (Rule Extraction)
This prompt guides the model to identify and summarize a distinguishing rule based on provided positive and negative examples. It emphasizes conciseness in the rule summary.
def visual_concept_prompt(m, n):
try:
if m < 0 or n < 0:
raise ValueError(f"Invalid input: m and n must be non-negative. Received m={m}, n={n}.")
if m > 0 and n > 0:
prompt = f"""
You are provided with {m + n} images: the first {m} samples are cat_2, the next {n} samples are cat_1. Analyze the common characteristics or patterns found in the cat_2 samples (positive samples: following 1 common rule) that distinctly separate them from the cat_1 samples (negative samples: it might not follow any possible rule).
Your task is to provide the rules that defines cat_2 samples. At the end, write "summary" of the rule identified in less than 20 words.
Ensure that the output is clear, well-formatted, and free of unnecessary explanations. Omit the ‘‘‘ tags at the beginning and end of the page.
if n == 0:
prompt = f"""
You are provided with {m} images: {m} samples are cat_2. Analyze the common characteristics or patterns found in the cat_2 samples (positive samples: following 1 common rule) that distinctly separate them from negative samples which might not follow any possible rule.
Your task is to provide the rules that defines cat_2 samples. At the end, write "summary" of the rule identified in less than 20 words.
Ensure that the output is clear, well-formatted, and free of unnecessary explanations. Omit the ‘‘‘ tags at the beginning and end of the page.
return prompt
except ValueError as e:
print(f"Error: {e}")
raise
A.5.2.2 Second-Stage Prompt (Rule Application)
This prompt presents the previously extracted rule summary and a query image, prompting the model to classify the image based on the rule. It reinforces the Bongard problem context and requests a structured output.
# Define the visual analysis prompt
def visual_concept_test_prompt(m, n, summary):
return f"""
[…]
Your task is to ponder over the rule and provide your conclusion for the ‘query image‘ if it can be categorized as either "cat_1" or "cat_2".
Ensure that the output is clear, well-formatted, and free of unnecessary explanations.
Omit the ‘‘‘ tags at the beginning and end of the page. The format of your output should be as follows:
- **Analysis**: (Your analysis here)
- **Rule**: (The distinguishing rule here)
- **Query Image**: (Query image details)
- **Conclusion**: (cat_1 or cat_2)
A.5.3 Componential Analysis
The Componential Analysis paradigm also uses a two-stage prompting strategy. The first stage generates detailed image descriptions, while the second stage derives a rule from these descriptions and applies it to a query image. The specific prompts for each stage are presented below.
A.5.3.1 First-Stage Prompt (Image Description Generation)
This prompt instructs the model to generate a comprehensive, hierarchical description of a given image in JSON format. It guides the model to cover various aspects of the image, from scene and objects to activities and contextual elements, facilitating detailed comparative analysis in the subsequent stage.
# Define the visual analysis prompt
def visual_concept_prompt():
Generates a visual analysis prompt.
Args:
Returns:
str: The formatted prompt string.
return """
Carefully examine the provided image and identify all possible visual elements, organizing them into a detailed hierarchical structure. Start with broad categories and progress to more specific subcategories. This should cover everything visible in the image, ensuring no detail is overlooked. Structure your findings in a JSON format to enable easy comparison and synthesis of data from other images. This will help discern patterns, contexts, and rules valuable for identifying or understanding query images.
Your hierarchy might encompass the following elements:
1. **Scene/Environment**: Description of the overall setting depicted, such as urban, natural, indoor, or outdoor scenes.
2. **Objects**: Define distinct items or entities present in the scene.
- **Living Beings**: Animals, humans, or other biological entities.
- Species or classification (e.g., dog, bird, human).
- Characteristics (e.g., color, posture, movement).
- **Inanimate Objects**: Both synthetic and natural elements.
- Categories (e.g., vehicle, building, trees).
- Properties (e.g., color, size, material, shape).
3. **Activities**: Observable actions or interactions involving any objects or beings.
- Specific descriptions of actions (e.g., walking, flying).
- Participants involved in these actions.
4. **Contextual Elements**: Environmental conditions and time markers, such as time of day or weather.
- Detailed characteristics (e.g., cloudy, night, winter).
5. **Visual Patterns**: Prominent colors, textures, and patterns that are visually significant.
6. **Emotional Undertones**: Any emotional presence or expressions evident in the image.
7. **Textual Information**: Any visible text within the image, including what it says and its visual style.
8. **Summary**: A concise narrative summarizing the overall content and context of the image.
Ensure that every aspect from the image is represented under these categories. The information should be presented in the following JSON format:
"Scene": {
"Description": "..."
"Objects": {
"Living Beings": [...],
"Inanimate Objects": [...]
"Activities": [...],
"Contextual Elements": {
"Time of Day": "...",
"Weather": "..."
"Visual Patterns": {
"Dominant Colors": [...],
"Textures": [...]
"Emotional Undertones": "..."
"Textual Information": "..."
"Summary": "..."
Ensure that the JSON output is clear, well-formatted, and free of unnecessary explanations. Omit the ‘‘‘json tags at the beginning and end of the page.
A.5.3.2 Second-Stage Prompt (Rule Derivation Instruction)
This prompt guides the model to analyze the JSON descriptions generated in the first stage, derive a distinguishing rule, and apply it to classify a query image. It emphasizes the use of the provided JSON format and requests a structured output.
def user_eval_prompt(all_image_specs, m, n):
return f"""
We are working with the Bongard dataset, which contains {m} images in cat_2 (positive samples) and {n} images in cat_1 (negative samples). These categories are defined as follows:
- Cat_2: Positive samples that follow a single common rule.
- Cat_1: Negative samples that may not follow any specific rule.
The image descriptions for the positive samples, negative samples, and the test image are provided in JSON format. Analyze the common patterns or characteristics in the cat_2 samples that distinguish them from cat_1 samples.
Your task is to:
1. Derive the rule that defines the cat_2 samples.
2. Apply this rule to categorize the test image.
Here are the image descriptions:
### Positive Samples (cat_2):
{all_image_specs[:m]}
### Negative Samples (cat_1):
{all_image_specs[m:m+n]}
### Test Image:
{all_image_specs[-1]}
Provide your output in the following format:
- **Analysis**: (Your analysis here)
- **Rule**: (The distinguishing rule here)
- **Test Image**: (Test image details)
- **Conclusion**: (cat_1 or cat_2)
Appendix A.6 Results
This appendix provides supplementary information and detailed results for the experiments described in the main paper. We present additional performance breakdowns and analyses to offer a more comprehensive view of the models’ capabilities on the Bongard OpenWorld dataset.
A.6.1 Componential Analysis Results by Commonsense Category
Table A.3 presents a detailed breakdown of the overall accuracy achieved by GPT-4o and Gemini 2.0 on the Bongard OpenWorld test set using the Componential Analysis paradigm. The results are categorized by Commonsense ID, providing insights into model performance across different types of commonsense reasoning challenges. This granular view allows for a more nuanced understanding of the models’ strengths and weaknesses in handling various aspects of commonsense knowledge.
ID Concept Category GPT 4o Gemini 2.0
0 Anything else 92.88 94.23
1 Human-Object Interaction (HOI) 86.67 92.86
2 Taste / Nutrition / Food 100.00 85.71
3 Color / Material / Shape 88.89 91.67
4 Functionality / Status / Affordance 100.00 100.00
5 And / Or / Not 90.00 80.00
6 Factual Knowledge 90.00 90.00
7 Meta Class 100.00 100.00
8 Relationship 100.00 100.00
9 Unusual Observations 92.31 92.31
Table A.3: Overall accuracy of GPT-4o and Gemini 2.0 on the Bongard OpenWorld test set using Componential Analysis, broken down by Commonsense ID.
A.6.2 Semantic Similarity in Deductive Rule Learning
Table A.4 presents the mean and standard deviation of the semantic similarity scores between the descriptions of the query images and the rules derived by the models during the Deductive Rule Learning experiments. These scores provide a measure of how well the extracted rules capture the semantic content of the images and the underlying Bongard problem. Analyzing these similarities helps to understand the relationship between the models’ ability to articulate rules and their overall performance on the task.
Gemini 2.0 GPT-4o
Category Mean Std Mean Std
Positive 0.915 0.02 0.9022 0.02
Negative 0.868 0.02 0.8659 0.02
Table A.4: Semantic Similarity between query image descriptions and identified Rules derived from Deductive Rule Learning
A.6.3 Impact of Chain-of-Thought Prompting
This section investigates the influence of Chain-of-Thought (CoT) prompting on VLM performance in the Direct Visual Rule Learning paradigm. CoT prompting, which encourages models to generate intermediate reasoning steps, has demonstrated significant improvements in complex reasoning tasks Wei et al. (2022b). We explore whether this approach is similarly beneficial in the visual domain, particularly for Bongard problems that demand a combination of perception and reasoning.
To assess the impact of CoT, we compare two prompting strategies: one that explicitly guides the model through a chain-of-thought by requesting a structured output (analysis, rule, query image description, and classification), and another that simply asks for the final classification.
Our analysis reveals a striking difference in performance between the two strategies. Without CoT prompting, the accuracy on the Direct Visual Rule Learning task drops significantly, from 80% to 61.60%. This decline is particularly pronounced for negative samples, where the accuracy falls from 66.40% to 39.20%, suggesting that the model struggles to identify instances where the rule does not apply when not explicitly prompted for a detailed reasoning process. This finding underscores the crucial role of CoT in guiding the model’s reasoning and enabling it to effectively navigate the complexities of visual reasoning tasks. These results are summarized in Table A.5.
Model Accuracy (%)
Overall neg pos
Without CoT 61.60 39.20 84.00
With CoT 80 66.40 93.60
Table A.5: Impact of CoT Prompting on Direct Visual Rule Learning when evaluated using GPT-4o model
The substantial impact of CoT observed in our experiments highlights its importance as a technique for eliciting more robust and reliable visual reasoning in VLMs. This suggests that incorporating CoT-like mechanisms into the model architecture or training process could lead to significant improvements in performance on complex visual tasks. Furthermore, our findings suggest that future research should explore the application of CoT prompting to other visual reasoning tasks and investigate the optimal strategies for integrating CoT into different VLM architectures.
A.6.4 Error Analysis Examples
This section provides detailed examples of Bongard OpenWorld test cases where both GPT-4o and Gemini 2.0 made incorrect classifications using the Componential Analysis paradigm. These examples illustrate common challenges faced by the models, offering insights into their limitations in visual reasoning and commonsense understanding. The analysis focuses on cases where both models failed, highlighting shared weaknesses. The “Reason” column explains the likely cause of misclassification based on the GPT-4o output (longer analysis, short rule, and test image description). These captions are reproduced from Wu et al. (2024).
Table A.6: Examples of Bongard OpenWorld Test Cases with Captions (Reproduced from Wu et al. (2024)) and Reasoning for Incorrect Classifications for images classified incorrectly by componential analysis by both by GPT-4o and Gemini 2.0. Human error analysis is based on the GPT-4o textual output: a longer analysis, short rule and the test image description.
SL NO Test ID Caption Reason
1 0021_neg_0 Cars on the city streets at night Weak reasoning (by similarity): a specific object in the rule (vehicles) is missing in the test image (impressionist painting), although deducible.
2 0014_neg_0 A person playing a guitar. Rule misses specific object (guitar) while being too general. The object is present in longer analysis.
3 0033_neg_0 A bicycle is placed in the corner Rule and test image analysis miss a specific property (in a corner / standing against a wall with ground visible). The test image is a collage of drawings on a wooden background.
4 0037_neg_0 The girl has long and thin braids on her head. Rule and test image analysis misses a specific property (specifically long and thin braids), being too general.
5 0076_pos_0 Various kinds of rings Rule misses a specific object (ring), being too general. Instead, rule focuses on a specific property (intricate design), which is missing in the test image and its description.
6 0076_neg_0 Various kinds of rings Rule misses a specific object (ring), being too general.
7 0082_neg_0 Live coral on the sea floor. Weak reasoning (by similarity): rule finds specific object (coral), but test image analysis does not.
8 0084_neg_0 A wooden fence surrounding a grassy field. Rule misses a specific object (grass), instead requiring greenery (grass or trees). The test image is a collage of natural images with greenery, no grass or trees are directly visible.
9 0112_neg_0 A wooden floor in the living room. Rule misses two specific objects (living room and a wooden floor), while focusing on just one object (wooden) and generalizing too much (indoor).
10 0117_neg_0 Colorful ribbons. Rule misses a specific object (colorful ribbons), generalizing too much.
11 0122_neg_0 A satellite view of Earth. Rule misses a specific viewpoint (top-down satellite view), using a too general (aerial view) instead.
12 0136_pos_0 Spectator seats view in the stadium. Weak reasoning: rule says “often active sports or spectators”, none of which are present in the test image or description. This lack is used as grounds for misclassification.
13 0213_neg_0 Checkerboard pattern fabrics Rule misses a specific object (fabric), the specific versions of which are present in analysis.
14 0234_neg_0 A beautiful stone sculpture Rule misses the correct property (tall), using instead another (prominent) of an obelisk. Test image features a prominent, but not a tall obelisk.
15 0247_pos_0 Small river filled with reeds Rule misses a specific object (reeds) present in analysis in a longer list, while prohibiting wrong objects (industrial presence), which are distantly present in test image.